OpenStudy (anonymous):

For the following equation, find two integers between -10 and 10 that make the equation true: -2x^2-2x=-4

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

How do I solve for x in this equation?

OpenStudy (anonymous):

well use the quadratic formula or if you want complete the square

OpenStudy (anonymous):

Ahhh that's right. Thank you

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

dont forget to set your equation in standard form. -2x^2-2x=-4 -2x^2-2x+4=0

OpenStudy (anonymous):

So once if I"m correct, 1 and -2 are integers that make the equation true?

OpenStudy (anonymous):

How would I find two integers betwen -10 and 10 that make the equation false?

OpenStudy (anonymous):

any number but -2 and 1 would be false :P

OpenStudy (anonymous):

any number betwen -10 and 10 not -2 or 1

OpenStudy (anonymous):

Oh of course, thank you again kind friend

OpenStudy (anonymous):

:)